# Revealing the Inherent Instructability of Pre-Trained Language Models

### v2

@@ Abstract
Instruction tuning—supervised fine-tuning using instruction-response pairs—is a key step in making pre-trained large language models (LLMs) instructable. Meanwhile, LLMs perform multitask learning during their pre-training, acquiring extensive knowledge and capabilities. We hypothesize that the pre-training stage can enable them to develop the ability to comprehend and address instructions. To verify this, we propose Response Tuning (RT), which removes the instruction and its corresponding mapping to the response from instruction tuning. Instead, it focuses solely on establishing the response distribution. Our experiments demonstrate that RT models, trained only on responses, can effectively respond to a wide range of instructions and exhibit helpfulness approaching that of their instruction-tuned counterparts. In addition, we observe that the models can recognize and reject unsafe queries after learning the refusal conditions from training responses. Furthermore, we demonstrate that these observations also hold in an in-context learning setting. These findings support our hypothesis, highlighting the extensive inherent capabilities of pre-trained LLMs. [^1] Our codes are available at https://github.com/seokhyunan/response-tuning.
@@
Seokhyun An1   Minji Kim2thanks: Work was partially conducted at UNIST.   Hyounghun Kim2footnotemark: 1Ulsan National Institute of Science and Technology (UNIST) 2Pohang University of Science and Technology (POSTECH) seokhyun@unist.ac.kr   {mzkim, imhkim}@postech.ac.kr
@@ 1 Introduction
Figure 1: Comparison of IT and RT. In both methods, the loss is computed exclusively on the response tokens. Unlike IT, RT omits the instruction and its corresponding mapping to the response, focusing solely on learning the response distribution.
@@
We hypothesize that LLMs can acquire the ability to process instructions—that is, to comprehend and address them—during pre-training, in addition to the extensive knowledge required to perform specific tasks (Radford et al., 2019; Brown et al., 2020; Zhou et al., 2024). For instance, they may implicitly learn instruction-response dynamics from Q&A threads on Stack Exchange—a common source in typical pre-training datasets (Elazar et al., 2024). Under this assumption, the models may be able to respond appropriately to instructions once their response distribution is established. Previous studies have reported similar phenomena in canonical NLP tasks, demonstrating that supervision from the output space is crucial for performing such tasks (Min et al., 2022; Kung and Peng, 2023).
@@
To test our hypothesis, we propose Response Tuning (RT), which omits the instruction and its associated instruction-response mapping from IT (see Figure 1). This omission prevents the model from learning to produce responses according to instructions; rather, it focuses on establishing the response distribution.
@@
We evaluate whether RT models can understand and respond appropriately to diverse instructions. In our experiments, we consider four recent LLMs and three different datasets: Llama-3.1-8B (Dubey et al., 2024), Gemma-2-2B and Gemma-2-9B (Riviere et al., 2024), and Mistral-7B-v0.3 (Jiang et al., 2023), utilizing only the responses from Alpaca (Taori et al., 2023), Dolly (Conover et al., 2023), and LIMA (Zhou et al., 2024). Our human and automatic evaluations, based on test instructions derived from multiple sources (Li et al., 2023; Lin et al., 2024), show that RT models can appropriately respond to a wide range of instructions and exhibit helpfulness close to that of their IT counterparts. These findings show that establishing a response distribution alone can yield instruction-following models, suggesting that the ability to process instructions is acquired during pre-training.
@@
Furthermore, we find that RT models can identify and reject unsafe requests by invoking learned refusal conditions. Specifically, we incorporate explanatory refusals for unsafe queries—responses that decline to fulfill requests while specifying the underlying safety policy—into the training data. When tested with diverse unsafe instructions derived from multiple safety evaluation datasets (Zou et al., 2023; Huang et al., 2024; Mazeika et al., 2024; Röttger et al., 2024), these models exhibit refusal rates approaching those of IT counterparts trained with paired examples. These results also support our hypothesis that pre-training enables the model to develop the ability to process instructions.
@@
Finally, we show that our observations hold in an in-context learning setting. We demonstrate that in-context learning with response-only demonstrations can enable the models to appropriately handle instructions, further supporting our hypothesis.
In summary:
@@
    Our extensive evaluations show that RT models—trained solely on responses—can effectively respond to a wide range of instructions, indicating that the ability to process instructions is acquired during pre-training.
@@
    Furthermore, we observe that they can recognize and reject unsafe requests by comprehending instructions and invoking learned refusal conditions. These observations are consistently replicated in an in-context learning setting, further supporting our hypothesis.
@@ 2 Related Work
Instruction tuning. Instruction Tuning (IT) is a supervised fine-tuning process using instruction-response paired data, where the model is trained to produce responses according to the instructions. Recent studies have shifted their focus from cross-task generalization in canonical NLP tasks (Weller et al., 2020; Mishra et al., 2022; Wei et al., 2022a; Sanh et al., 2022) to generalization for unseen user instructions. Notable contributions include synthetic data generation frameworks (Wang et al., 2023b; Honovich et al., 2023; Ding et al., 2023; Xu et al., 2024), human-involved conversation collection methods (Conover et al., 2023; Köpf et al., 2023; Chiang et al., 2023; Zhou et al., 2024; Zhao et al., 2024; Zheng et al., 2024). Recently, a line of research has explored the superficiality of the IT stage, questioning the necessity of large-scale data (Zhou et al., 2024; Chen et al., 2024; Liu et al., 2024) or parameter updates (Lin et al., 2024). However, what the models learn from IT remains unclear (Kung and Peng, 2023; Ghosh et al., 2024), and ablation studies exploring these aspects are limited.
Learning from input-output pairs. Min et al. (2022) highlights the crucial role of label space information in the in-context learning of canonical NLP tasks (e.g., classification). Kung and Peng (2023) examines the role of task definitions in the prompts of the IT dataset and demonstrates that models trained with misleading task definitions or only on the label space exhibit similar task generalization performance. However, these studies focus on generalization across canonical tasks with a finite label space, rather than on how models handle natural language instructions with an open-ended response space.
Our approach. To address these gaps, we investigate whether RT models—which have not been explicitly trained to process instructions—can understand and respond to them appropriately. A concurrent study by Hewitt et al. (2024) also examines whether training only on responses can yield instruction-following. In our work, we carefully investigate the inherent ability to process instructions using a multi-pronged evaluation, which includes refusal tests for unsafe queries and assessments of in-context response learning.
@@ 3 Response Tuning (RT)
We propose Response Tuning (RT) to verify our hypothesis that the ability to process instructions can be developed in the pre-training stage.
@@
where $\text{where }\delta_{i}$ =$=\begin{cases}1&\text{if }t_{i}\in\mathbf{R}\\ 0&\text{otherwise }\end{cases}$
@@
Inference. We input a sequence that starts with the instruction delimiter (<|user|>), followed by the user’s instruction, and then the response delimiter (<|assistant|>). The model then generates the assistant’s response after the <|assistant|> token. Although RT does not involve explicitly learning how to process instructions, we find that the models can effectively leverage their inherent capabilities to do so. We demonstrate this in the following experimental sections.
@@ 4.1 Experimental Setup
Training dataset. We use three different IT datasets, from which we only use the response subsets: [^2] Theoretically, RT can also be performed on general texts that lack paired instructions. We explore this approach in Appendix F.
@@
Figure 4: Pairwise evaluation results for Llama-3.1-8B-based RT models. Human evaluators and the GPT-4 judge are asked to choose the more helpful response between RT and IT models for the same instruction. The results indicate that RT models exhibit preferences similar to their IT counterparts. The results for the other models are provided in Tables 9 and 10.
@@
Untuned 69.04 33.8 61.09 74.42 69.9 81.28 64.92
Table 1: Core capabilities evaluation results for IT and RT models. The results indicate that RT models largely retain their core capabilities and exhibit performance similar to that of IT models across all benchmarks. See Table 9 for results on the other models.
@@
Figure 6: Refusal evaluation results for RT and IT models trained on datasets that include refusal cases. The results indicate that RT models can refuse unsafe queries and achieve refusal rates approaching those of IT models trained on paired data. We find no substantial differences in false refusal rates between the two models. See Tables 11 and 12 for the results of the other models.
Instructabiltiy evaluation. To evaluate whether RT models can handle user instructions, we assess their responses’ acceptability and quality. Additionally, we conduct pairwise assessments to measure their relative helpfulness compared to their IT counterparts.
4.2 Results
RT models can appropriately respond to instructions. Figures 23 and 4 depict the human acceptability, response quality, and pairwise assessment results, respectively. The independent assessment results show that RT models—trained solely on responses—can generate appropriate answers to diverse instructions. A majority of the responses generated by RT models are rated as Acceptable, with many achieving the Excellent rating. They perform close to their IT counterparts in response quality evaluations, achieving similar scores across all metrics of the JustEval benchmark. In contrast, we observe that the base models rarely produce valid responses due to the absence of an established response distribution.
Additionally, the pairwise evaluation results indicate that while RT models can produce appropriate responses, their outputs are slightly less preferred compared to their IT counterparts. Our manual investigation of cases where IT model outputs are preferred over those of RT models finds that IT models often generate responses in a format better suited to the given instructions. These suggest that explicit instruction-response mapping helps models select a response format that is more appropriate from the learned distribution of responses. The examples of the model outputs are available in Appendix D. We further investigate how the training response distribution affects the model’s output in Appendices F and G.
In summary, our instructability evaluation results indicate that RT models—which are not explicitly trained to process instructions but rather learn solely the response distribution—can still provide helpful responses to diverse instructions. This finding suggests that instruction-processing ability is acquired during pre-training. [^4] We discuss the potential contamination concern regarding the IT datasets in Appendix H.
RT models mostly retain their core capabilities. Table 1 presents the core capabilities evaluation results. We find that RT models largely retain their core capabilities, with no notable degradation beyond the typical performance decline associated with fine-tuning (Wang et al., 2023a; Fu et al., 2024). Furthermore, across all tested tasks, RT and IT models exhibit similar performance, showing no substantial differences. These results suggest that RT models possess a knowledge base similar to that of IT models, which is required to perform the instructed tasks.
5 Rejecting Unsafe Instructions
In this section, we further investigate whether RT models can comprehend instructions and decide how to handle them. To test this, we introduce refusals for unsafe queries only through responses—thereby avoiding direct training for refusals via instruction-response mappings—and then evaluate whether the models can recognize and reject unsafe instructions.
5.1 Experimental Setup
@@ 6 In-Context Response Learning
Experimental setup. We remove instructions and the associated instruction-response mappings from URIAL (Lin et al., 2024), which consists of 4 instruction-response pairs including one pair of unsafe instruction and refusal. We refer to this new version as URIAL-R. We then evaluate it using two different base LLMs, Llama-3.1-8B, and Gemma-2-9B, with the JustEval benchmark. We employ greedy decoding with a maximum generation length of 2,048 tokens. We also evaluate the zero-shot template prompting baseline (Lin et al., 2024) to further investigate the effect of learning the response distribution. The prompts and details of the setup can be found in Appendix A.
Results. Figure 7 presents the evaluation results. The results show that using response demonstrations alone enables the base models to effectively handle both benign and unsafe instructions. Across all metrics, the scores of the model prompted with URIAL-R are similar to those of the model prompted with the original URIAL, which includes instructions. Additionally, while the zero-shot prompting baseline generates outputs relevant to the instructions, URIAL-R—which includes demonstrations of coherent responses—substantially outperforms it. These results further reinforce our earlier conclusion that the ability to handle instructions is inherent in pre-trained LLMs, and that establishing an adequate output distribution helps the models use them effectively.

### v3

@@ Abstract
Instruction tuning—supervised fine-tuning using instruction-response pairs—is a key step in making pre-trained large language models (LLMs) instructable. Meanwhile, LLMs perform multitask learning during their pre-training, acquiring extensive knowledge and capabilities. We hypothesize that the pre-training stage can enable them to develop the ability to comprehend and address instructions. To verify this, we propose Response Tuning (RT), which removes the instruction and its corresponding mapping to the response from instruction tuning. Instead, it focuses solely on establishing a response distribution. Our experiments demonstrate that RT models, trained only on responses, can effectively respond to a wide range of instructions akin to their instruction-tuned counterparts. In addition, we observe that the models can recognize and reject unsafe queries after learning a safety policy only from the response data. Furthermore, we find that these observations extend to an in-context learning setting. These findings support our hypothesis, highlighting the extensive inherent capabilities of pre-trained LLMs. [^1] Our codes are available at https://github.com/seokhyunan/response-tuning.
@@
Seokhyun An1   Minji Kim2thanks: Part of this work was conducted at UNIST.   Hyounghun Kim2,3footnotemark: 1Department of Computer Science and Engineering, UNIST 2Graduate School of Artificial Intelligence, POSTECH 3Department of Computer Science and Engineering, POSTECH seokhyun@unist.ac.kr   {mzkim, h.kim}@postech.ac.kr
@@ 1 Introduction
Figure 1: Comparison of IT and RT. In both methods, the loss is computed exclusively on the response tokens. Unlike IT, RT omits the instruction and its corresponding mapping to the response, focusing solely on learning a response distribution.
@@
We hypothesize that LLMs can acquire the ability to process instructions—that is, to comprehend and address them—during pre-training, in addition to the extensive knowledge required to perform specific tasks (Radford et al., 2019; Brown et al., 2020; Zhou et al., 2024). For instance, they may implicitly learn instruction-response dynamics from Q&A threads on Stack Exchange—a common source in typical pre-training datasets (Elazar et al., 2024). Under this assumption, the models may be able to respond appropriately to instructions once their response distributions are established. Previous studies have reported similar phenomena in canonical NLP tasks, demonstrating that supervision from the output space is crucial for performing such tasks (Min et al., 2022; Kung and Peng, 2023).
@@
To test our hypothesis, we propose Response Tuning (RT), which omits the instruction and its associated instruction-response mapping from IT (see Figure 1). This omission prevents the model from learning to produce responses according to instructions; rather, it focuses on establishing a response distribution.
@@
We evaluate whether RT models can understand and respond appropriately to diverse instructions. In our experiments, we consider four recent LLMs and three different datasets: Llama-3.1-8B (Dubey et al., 2024), Gemma-2-2B and Gemma-2-9B (Riviere et al., 2024), and Mistral-7B-v0.3 (Jiang et al., 2023), utilizing only the responses from Alpaca (Taori et al., 2023), Dolly (Conover et al., 2023), and LIMA (Zhou et al., 2024). Our human and automatic evaluations, based on test instructions derived from multiple sources (Li et al., 2023; Lin et al., 2024), show that RT models can appropriately respond to a wide range of instructions and exhibit helpfulness close to that of their IT counterparts. These findings show that establishing a response distribution alone can yield instruction-following models, suggesting that pre-training allows the models to learn how to process instructions.
@@
Furthermore, we find that RT models can identify and reject unsafe queries. Specifically, we incorporate explanatory refusals for unsafe queries—responses that decline to fulfill requests while specifying the underlying safety policy—into the training data. When tested with diverse unsafe instructions derived from multiple safety evaluation datasets (Zou et al., 2023; Huang et al., 2024; Mazeika et al., 2024; Röttger et al., 2024), these models invoke their learned safety rules to the unsafe queries, exhibiting refusal rates approaching those of IT counterparts trained with paired examples. These results also support our hypothesis that pre-training enables the model to develop the ability to process instructions.
@@
Finally, we extend our investigation to an in-context learning setting to test if our prior observations hold. We find that response-only demonstrations can enable the models to appropriately handle instructions, further supporting the hypothesis.
Overall, our findings provide a deeper understanding of how LLMs become instructable agents and suggest the potential of extensive inherent capabilities acquired during pre-training. In summary:
@@
    Our extensive evaluations show that RT models—trained solely on responses—can effectively respond to a wide range of instructions. Furthermore, we observe that they can recognize and correspondingly reject unsafe requests by invoking a learned safety policy. These results indicate that LLMs learn how to process instructions during pre-training.
@@
    These findings extend to an in-context learning setting, further supporting our hypothesis.
@@ 2 Related Work
Instruction tuning. Instruction Tuning (IT) is a supervised fine-tuning process using instruction-response paired data, where the model is trained to produce responses according to the instructions. Recent studies have shifted their focus from cross-task generalization in canonical NLP tasks (Weller et al., 2020; Mishra et al., 2022; Wei et al., 2022a; Sanh et al., 2022) to generalization for unseen user instructions. Notable contributions include synthetic data generation frameworks (Wang et al., 2023b; Honovich et al., 2023; Ding et al., 2023; Xu et al., 2024), human-involved conversation collection methods (Conover et al., 2023; Köpf et al., 2023; Chiang et al., 2023; Zhou et al., 2024; Zhao et al., 2024; Zheng et al., 2024). Recently, a line of research has explored the superficiality of the IT stage, questioning the necessity of large-scale data (Zhou et al., 2024; Chen et al., 2024; Liu et al., 2024) or parameter updates (Lin et al., 2024). However, what the models learn from IT and how they become instructable agents remains unclear (Kung and Peng, 2023; Ghosh et al., 2024).
Learning from input-output pairs. Min et al. (2022) highlights the crucial role of label space information in in-context learning of canonical NLP tasks (e.g., classification). Kung and Peng (2023) examines the role of task definitions in the prompts of the IT dataset and demonstrates that models trained with misleading task definitions or only on the label space exhibit similar task generalization performance. However, these studies focus on generalization across canonical tasks with a finite label space, rather than on how models handle natural language instructions with an open-ended response space.
Our approach. To address these gaps, we employ RT to investigate how well LLMs can process given instructions, without additional fine-tuning on the paired data, but with an established response distribution. A concurrent study by Hewitt et al. (2024) examines whether training only on responses can yield instruction-following. In this work, we focus on the inherent ability to process instructions, employing various evaluation methods (including tests for unsafe query identification) across both fine-tuning and in-context learning settings to carefully assess them.
@@ 3 Response Tuning (RT)
We propose Response Tuning (RT), a simple ablation of IT. It prevents further learning of instruction–response mappings by omitting the instructions, thereby allowing us to investigate the model’s capability to process instructions acquired during pre-training.
@@
$\text{where }\delta_{i}$ $=\begin{cases}1&\text{if }t_{i}\in\mathbf{R}\\ 0&\text{otherwise }\end{cases}$
@@
Inference. We input a sequence that starts with the instruction delimiter (<|user|>), followed by the user’s instruction, and then the response delimiter (<|assistant|>). The model then generates the assistant’s response after the <|assistant|> token. Although RT provides no explicit supervision for instruction-processing, we find that the models can effectively leverage their inherent capabilities to do so. We demonstrate this in the following experimental sections.
@@ 4.1 Experimental Setup
Training dataset. We use three different IT datasets, from which we only use the response subsets: [^2] Theoretically, RT can be performed on general texts that lack paired instructions. We explore this approach in Appendix F.
@@
Figure 4: Pairwise evaluation results for Llama-3.1-8B-based RT models. Human evaluators and the GPT-4 judge are asked to choose the more helpful response between RT and IT models for the same instruction. The results indicate that RT models exhibit preferences similar to their IT counterparts. The results for the other models are provided in Table 9 and 10.
@@
Untuned 69.04 33.80 61.09 74.42 69.90 81.28 64.92
Table 1: Core capabilities evaluation results for RT and IT models. The results indicate that RT models largely retain their core capabilities and exhibit performance similar to that of IT models across all benchmarks. See Table 12 for results on the other models.
Instructability evaluation. To evaluate whether RT models can handle user instructions, we assess their responses’ acceptability and quality. Additionally, we conduct pairwise assessments to measure their relative helpfulness compared to their IT counterparts.
Query
Think of topics that are most common in classic interview questions for a job in computer science.
[…]
1. Data Structures and Algorithms: common algorithms such as sorting and searching, time and space […]
2. Programming Languages and Paradigms: familiarity with specific programming languages […]
[…]
Some common topics covered in a computer science job interview include data structures and algorithms, object-oriented design and programming […]
Table 2: Outputs from Llama-3.1-8B RT and IT models trained on Alpaca. The auto-evaluator preferred the output from the IT model that provides an enumeration-style response.
Model InFoBench
Llama-3.1-8B IT 0.77
++Alpaca RT 0.69
Gemma-2-9B IT 0.79
++Alpaca RT 0.74
Table 3: Average DRFR for RT and IT models. The models are evaluated on 500 decomposable test instructions from InFoBench (Qin et al., 2024). The results indicate that models trained with paired examples more sensitively capture prompt requirements.
Query
Generate a list of 5 open-ended questions about thermodynamics suitable for a high school physics exam, with each question containing at least two keywords about thermodynamics, which should be provided in brackets and separated by a semicolon after the question.
[…]
1. How do changes in temperature and pressure affect the thermodynamic processes of an ideal gas [temperature; pressure]?
2. Can you explain the concept of heat transfer and the factors that affect it [heat; transfer]? […]
[…]
1. [Heat, Temperature, Expansion] How does the increase in temperature of an object affect its expansion, and how is this phenomenon explained thermodynamically? (Please provide a brief explanation and an example.)
2. [Work, Heat, Energy] Can heat ever be considered a form of work, and if so, under what thermodynamic conditions does this apply? (Please provide a brief explanation and an example.) […]
Table 4: Outputs from Llama-3.1-8B RT and IT models for a decomposable instruction. Both models produce largely valid responses; however, the RT model misses some prompt requirements, whereas the IT model captures them exactly.
@@ 2. Can you explain the concept of heat transfer and the factors that affect it [heat; transfer]? […]
4.2 Results
RT models can appropriately respond to instructions. Figure 23 and 4 depict the human acceptability, response quality, and pairwise assessment results, respectively. The independent assessment results show that RT models—trained solely on responses—can generate appropriate responses to diverse instructions. A majority of the responses generated by RT models are rated as Acceptable, with many achieving the Excellent rating. They perform close to their IT counterparts in response quality evaluations, achieving similar scores across all metrics of the JustEval benchmark. Consistently, the pairwise evaluation results indicate that RT models can produce appropriate responses like IT models. In contrast, we observe that the base models rarely produce valid responses due to the absence of an established response distribution. [^4] We discuss the potential contamination concern regarding the IT datasets in Appendix H. These results suggest that pre-training allows the models to develop the instruction-processing ability. The examples of the model outputs are available in Appendix D. We further investigate how the training response distribution affects the model’s output in Appendix F and G.
Explicit instruction–response mappings improve the formatting of responses and task sensitivity. While RT models can produce appropriate responses, their outputs are slightly less preferred than those of their IT counterparts. In our manual inspection of such cases, we occasionally observe that IT models produce better-formatted responses for tasks that favor specific response formats (e.g., suggestions or brainstorming). For example, in Table 2, the IT model generates a numbered list that better matches the brainstorming task. This suggests that explicit instruction–response mapping helps models deliver responses that are better structured for the task.
@@ 4.2 Results
Figure 6: Refusal evaluation results for RT and IT models trained on datasets including refusal examples. The results indicate that RT models can identify and reject unsafe queries and achieve refusal rates approaching those of IT models trained on paired data. We find no substantial differences in false refusal rates between the two models. See Table 15 and 16 for the results of the other models.
@@
Additionally, we further quantify differences with InFoBench (Qin et al., 2024), which computes the Decomposed Requirement-Following Ratio (DRFR)—the proportion of individual prompt requirements that a model satisfies—on a set of decomposable instructions. The results in Table 3 show that, although RT models achieve DRFR approaching those of IT models, IT models capture prompt requirements more sensitively. For instance, in Table 4, the instruction explicitly requests keywords separated by semicolons following the questions. Both models produce largely valid responses, but the RT model separates the keywords with commas and places them before the questions, whereas the IT model captures the requirements exactly. This suggests that the explicit mapping helps the models more sensitively capture prompt requirements. The results for the other models are available in Appendix C.
RT models mostly retain their core capabilities. Table 1 presents the core capabilities evaluation results. We find that RT models exhibit similar performance to IT models and largely retain their core capabilities. These results suggest that RT models possess a knowledge base similar to that of IT models, which is required to perform the instructed tasks.
5 Rejecting Unsafe Instructions
In this section, we investigate whether RT models can understand instructions and correspondingly decide how to handle them. To test this, we incorporate refusals for unsafe queries into RT training data and then evaluate whether they can recognize and reject unsafe instructions. Unlike IT, which utilizes explicit unsafe instruction-rejection mappings to allow models to identify them, RT omits such direct mappings.
5.1 Experimental Setup
@@ 6 In-context Response Learning
Experimental setup. We remove instructions and the associated instruction-response mappings from URIAL (Lin et al., 2024), which consists of 4 instruction-response pairs including one pair of unsafe instruction and refusal. We refer to this new version as URIAL-R. We then evaluate it using two different base LLMs, Llama-3.1-8B, and Gemma-2-9B, with the JustEval benchmark. We employ greedy decoding with a maximum generation length of 2,048 tokens. We also evaluate the zero-shot template prompting baseline (Lin et al., 2024) to further investigate the effect of learning a response distribution. The prompts and details of the setup can be found in Appendix A.
Results. Figure 7 presents the evaluation results. The results show that using response demonstrations alone enables the base models to effectively handle both benign and unsafe instructions. Across all metrics, the scores of the models prompted with URIAL-R are similar to those of the models prompted with URIAL, which includes instructions. Additionally, while the zero-shot prompting baseline generates outputs relevant to the instructions, URIAL-R—which includes demonstrations of coherent responses—substantially outperforms it. These results further reinforce our earlier conclusion that the ability to handle instructions is inherent in pre-trained LLMs and that establishing an adequate output distribution helps the models use them effectively.